OpenStudy (anonymous):

Let f(x)=|3x| . Compute the limit lim h→0 (f(0+h)-f(0))/(h)

OpenStudy (anonymous):

f(x)=|3x| f(x+h) = |3(x+h)| f(0) = 0 f(0+h) = |3(0+h)| f(0+h) = |3h| can you find this limit now???

OpenStudy (anonymous):

no, it's still saying it's incorrect. Thank you though

OpenStudy (anonymous):

\[\frac{ f(0+h)-f(0) }{ h }\] this is the format they give me

OpenStudy (anonymous):

lim h→0 (f(0+h)-f(0))/(h) = lim h→0 3h/h = lim h→0 3 = 3 ---> is your answer

OpenStudy (anonymous):

yes, that's what i entered and it's still saying it's incorrect

OpenStudy (anonymous):

so check your question may be it is wrong

OpenStudy (anonymous):

nope, that's exactly as the question reads. i've been working on this for 45 minutes and can't get it.

OpenStudy (anonymous):

o o i got it

OpenStudy (anonymous):

as f(x)=|3x| then f(0+h) = |3h| now f(0+h) = -3h for 3h<0 and f(0+h) = 3h for 3h>0

OpenStudy (anonymous):

I finally got it! thank you for all your help!

OpenStudy (anonymous):

goooood